# Permutation-Invariant Spectral Learning via Dyson Diffusion

### v1

@@ Abstract
Diffusion models are central to generative modeling and have been adapted to graphs by diffusing adjacency matrix representations. The challenge of having up to $n!$ such representations for graphs with $n$ nodes is only partially mitigated by using permutation-equivariant learning architectures. Despite their computational efficiency, existing graph diffusion models struggle to distinguish certain graph families, unless graph data are augmented with ad hoc features. This shortcoming stems from enforcing the inductive bias within the learning architecture. In this work, we leverage random matrix theory to analytically extract the spectral properties of the diffusion process, allowing us to push the inductive bias from the architecture into the dynamics. Building on this, we introduce the Dyson Diffusion Model, which employs Dyson’s Brownian Motion to capture the spectral dynamics of an Ornstein–Uhlenbeck process on the adjacency matrix while retaining all non-spectral information. We demonstrate that the Dyson Diffusion Model learns graph spectra accurately and outperforms existing graph diffusion models.
@@ 1 Introduction
Diffusion models are a key class of generative models based on noising data with Stochastic Differential Equations (SDEs) and learning their time reversal (Sohl-Dickstein et al., 2015; Song et al., 2021; Ho et al., 2020). They provide state-of-the-art results in many domains such as audio (Zhang et al., 2023) and vision (Croitoru et al., 2023). Generalizing diffusion models from Euclidean space to graphs offers promising applications in numerous areas, such as biology (Watson et al., 2023) or combinatorial optimization (Sun & Yang, 2023). However, while diffusing adjacency matrix representations is straightforward and popular (Niu et al., 2020; Jo et al., 2022; Vignac et al., 2022), this approach faces a major obstacle: Each graph with $n$ vertices has up to $n!$ representations as adjacency matrices. Therefore, if one aims to use a diffusion model on the space of matrices, one must learn $n!$ representations per graph. This is not feasible. Previous works tackled this problem by shifting the inductive bias of permutation invariance to the learning algorithm: If the neural network was permutation equivariant, training on one of the (up to $n!$ many) matrix representations would suffice. For example, Niu et al. (2020) and Jo et al. (2022) used message-passing graph neural networks (GNNs) while ConGress (Vignac et al., 2022) applied graph transformers. However, these learning architectures have a “blind spot” detailed below.
@@
Theoretical Limitations.  The “blind spot” arises from the limited ability of these models to solve Graph Isomorphism ($\mathrm{GI}$): determining if two graphs are structurally identical regardless of node labeling. While permutation equivariance ensures that the model produces consistent outputs for different representations of the same graph, it does not guarantee that the model can differentiate between two structurally different (non-isomorphic) graphs. The failure is a result of how these architectures aggregate information: permutation-symmetric operations – message passing in GNNs and self-attention in graph transformers – can collapse distinct graphs with similar neighborhoods to the same representation, treating them as identical. More formally, $\mathrm{GI}$ is a challenging problem in algorithm theory, and it remains unknown whether $\mathrm{GI}\in\mathsf{P}$ (Babai, 2016). Since a polynomial time (learning) algorithm perfectly distinguishing all graphs would prove $\mathrm{GI}\in\mathsf{P}$, contemporary (polynomial-time) graph learning algorithms must compromise on expressivity.
@@
Extracting Permutation-Invariant Information from Graph Diffusion. When diffusing an entire adjacency matrix, state-of-the-art work pushes the entire inductive bias into the learning algorithm (Niu et al., 2020; Jo et al., 2022) with possible data augmentation (Huang et al., 2022; Vignac et al., 2022; Xu et al., 2024). However, as we show below, this is neither necessary nor desirable (see the previous discussion on blind spots). Using techniques from random matrix theory, we show that an OU diffusion on the graph can be dissected into diffusion of the (permutation invariant) spectrum and diffusion of the (permutation-dependent) eigenvectors. Our method therefore allows us to learn the spectrum while preserving all remaining information. Moreover, since the spectrum is inherently permutation-invariant, we can parameterize the score using a much broader range of learning architectures, expanding the scope to architectures able to distinguish between graphs that are equivalent in the Weisfeiler-Leman (WL) sense (Morris et al., 2019; Xu et al., 2018). [^1] We give a theoretical discussion in form of the WL-equivalence class in Section 2 and demonstrate the challenge of those architectures empirically in Figure 2.
@@
Figure 1: Dyson Diffusion model and its application to graph spectra: A graph on $n$ vertices (a) has up to $n!$ representations as adjacency matrices (b). For an OU-driven diffusion on any adjacency matrix, the permutation-invariant spectrum (c) evolves according to the same SDE (Dyson-BM). An exemplary path of the $n$, non-intersecting, eigenvalues is shown. The marginals of the invariant density for the $\lambda_{k}$ are depicted on the far right. The DyDM diffusion model learns the score $s(\lambda,t)$ (highlighted in yellow) to generate spectra via the time-reversed SDE (5).
@@
Information in the Spectrum. The spectrum of a graph encodes key structural features including connectivity, expansion, and subgraph patterns. Moreover, non-isomorphic WL-equivalent graphs typically have distinct spectra (Huang & Yau, 2024). An idea, therefore, is to augment the graph data based on spectral information (Vignac et al., 2022; Xu et al., 2024). Our work is based on an entirely different method, exploiting analytical insight from random matrix theory to dissect the spectral from the remaining information, allowing to push the inductive bias from the architecture to the dynamics.
@@
Dyson’s Brownian Motion. For an OU process on the space of symmetric matrices, the eigenvalues follow a well-characterized stochastic differential equation (SDE), the so-called Dyson Brownian Motion (DBM), which is inherently permutation invariant. Thus – in contrast to Niu et al. (2020); Jo et al. (2022); Vignac et al. (2022) – the score of DBM can be parameterized with any (not necessarily permutation invariant) neural network. Moreover, contrary to Luo et al. (2024), the remaining information is not lost (see Theorem 3.2).
Contributions. The main contributions of this work are as follows.
  • We introduce the novel Dyson Diffusion Model (DyDM) in Section 3, which extracts the spectral dynamics from an OU-driven diffusion. DyDM allows to learn the spectra of graphs without the need of GNNs or graph transformers while preserving the remaining information of the graph and allowing us to compute an eigenvector SDE.
  • We demonstrate in Section 4 that DyDM is more effective than existing GNN-based and graph-transformer-based methods for learning graph spectra.
@@
  • We illustrate the struggle of GNN-based graph diffusion models in Figure 2 and Section 4.1.
@@
Beyond graphs, our framework applies generally to symmetric matrices where spectral information is often key. For instance, in statistics, it can encode the importance of principal components (James et al., 2023), while in dynamical systems, [^2] For instance, Markov jump dynamics in detailed balance systems where the generator is symmetric in the steady state basis, see Pavliotis (2014). it reflects the stability and timescales of linear operators.
@@
Notation. We work on the set of symmetric real matrices $\mathrm{Sym}(\mathbb{R}^{n\times n}):=\{A\in\mathbb{R}^{n\times n}:A^{T}=A\}$. The positive integers up to $n$ are denoted by $n]:=\{1,\ldots n\}$. We consider undirected graphs $G=(V,E,w)$ where $V$ is a finite set with edges $E\subseteq\{S\subseteq V:1\leq|S|\leq 2\}$ allowing for self-loops and weights $w:E\to\mathbb{R}$. The family of such graphs of size $n$ is $\mathcal{G}^{n}:=\{G=(V,E,w):|V|=n\}$. For a graph $G\in\mathcal{G}^{n}$ with $\mathbb{R}^{V}=\{f:V\to\mathbb{R}\}$ being the space of functions from $V$ to $\mathbb{R}$, we let $A\equiv A(G):\mathbb{R}^{V}\to\mathbb{R}^{V}$ be the adjacency operator, defined for $f\in\mathbb{R}^{V}$ and $v\in V$ as the weighted sum of $f$ applied to the neighbors of $v$ as
$Af(v)\coloneqq\sum_{\{u,v\}\in E}w\left(\{u,v\}\right)f(u).$
As the graphs are undirected, the adjacency operator is self-adjoint with respect to the standard inner product on $\mathbb{R}^{V}$. The operator $A$ therefore has $n$ real eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{n}$. We denote the ordered spectrum of the adjacency operator by
$\lambda(G)\coloneqq\lambda(A(G))\coloneqq\{(\lambda_{1},\lambda_{2},\ldots,\lambda_{n})\,:\,\lambda_{1}\geq\ldots\geq\lambda_{n}\}.$
Unless defined explicitly otherwise, we use $n\in\mathbb{N}$ for the (vertex) size of the graph, and $N\in\mathbb{N}$ for the number of samples. We denote by $s\sim\mathcal{N}(0,I_{d})$ that the $d$-dimensional random vector $s$ has multivariate normal distribution with 0 mean and unit covariance $I_{d}$.
@@ 2 Limitations of Existing Diffusion Models
We consider the following matrix-valued Ornstein–Uhlenbeck (OU) SDE starting from some data $M(0)\in\mathrm{Sym}(\mathbb{R}^{n\times n})$ given for $1\leq i\leq j\leq n$ by
$\mathrm{d}M_{ji}(t)=\mathrm{d}M_{ij}(t)=-\beta M_{ij}(t)\mathrm{d}t+D_{ij}\mathrm{d}B_{ij}(t),$ (1)
@@
with diffusion coefficient $D_{ij}\coloneqq\sqrt{(1+\delta_{ij})\alpha}$ for any constants $\alpha,\beta\in\mathbb{R}^{+}$, where $B_{ij}(t)=B_{ji}(t)$ are independent Brownian motions and $\delta_{ij}=1$ if and only if $i=j$, and 0 otherwise. We consider eq. (1) on the space of symmetric matrices to represent undirected graphs $G\in\mathcal{G}^{n}$.
Equation 1 is an entry-wise OU process that preserves the symmetry of the matrix. In standard diffusion models we run eq. 1 until some time $T>0$ from data samples. To fix notation, denote by $p_{t}$ the distribution of $M(t)$ induced by (1) for $t\geq 0$.
The time-reversal of the diffusion (1) over the time interval $[0,T]$ is a diffusion initialized by sampling $M(T)\sim p_{T}$ and satisfying for $1\leq i\leq j\leq n$
$\mathrm{d}M_{ij}(t)=-\left\{\beta M_{ij}(t)+D_{ij}^{2}\left[s(M(t),t)\right]_{ij}\right\}dt+D_{ij}\mathrm{d}\bar{B}_{ij}(t),$ (2)
@@
for independent Brownian motions $\bar{B}_{ij}(t)$ (Anderson, 1982; Song et al., 2021). In Equation 2, $s(M,t)$ represents the score matrix at time $t\in(0,T]$, i.e., $s(M,t)\coloneqq\gradient_{M}\log p_{t}(M)$. This score is intractable but, as the OU process has tractable Gaussian transition densities, we can obtain an estimate $s_{\theta}(M,t)$ of it by minimizing the denoising score matching loss (Song et al., 2021)
$L(\theta)=\operatorname{\mathbb{E}}_{t\sim\mathrm{Unif}[0,T],M(0)\sim p_{0},M(t)\sim p_{t\mid 0}(\cdot|M(0))}\left[\norm{s_{\theta}(M(t),t)-\gradient_{M(t)}\log p_{t}(M(t)|M(0))}^{2}_{2}\right].$ (3)
@@
Approximate samples of $p_{0}$ can then be obtained by simulating an approximation of Equation 2 obtained by sampling $M(T)$ from the Gaussian invariant distribution of (1), that is, $M_{ij}^{\rm inv}\sim\mathcal{N}(0,\alpha(1+\delta_{ij})/2\beta)$, and using $s_{\theta}(M,t)$ in place of $s(M,t)$.
@@
Challenges posed by graphs. When working with graphs, to obtain an adjacency matrix for a given weighted graph $G=(V,E,w)\in\mathcal{G}^{n}$, we need to fix an ordering $(v_{1},\ldots,v_{n})$ of all vertices. In fact, given the ordering $(v_{1},\ldots,v_{n})$, the associated matrix $M=(m_{ij})_{1\leq i,j\leq n}\in\mathrm{Sym}(\mathbb{R}^{n\times n})$ has entries $m_{ij}=w(\{v_{i},v_{j}\})$ if $\{v_{i},v_{j}\}\in E$ and 0 otherwise. The challenge is that the adjacency operator $A(G)$ for a given graph $G\in\mathcal{G}^{n}$ admits up to $n!$ distinct adjacency matrices. For instance, in Figure 1 we see four different matrix representations of the same graph.
Why should we enforce this inductive bias? One could think that the reason for using the inductive bias stems from wanting that graph generative models assign uniform probability to each of the (up to) $n!$ many representations. But this can be easily achieved by applying a random independent permutation to the output of the generative model.
@@
Instead, the challenge stems from the following problem: In diffusion models, we learn a function (i.e. the score) on a set of graphs, say [^3] Importantly, $\Omega$ refers to the set of graphs and not the set of (permutation-sensitive) adjacency matrices. $\Omega$, rather than the distribution directly. Learning on the adjacency representations would correspond to learning on $\Omega\times S_{n}$. We demonstrate on a toy example in Corollary K.1 that not leveraging the inductive bias, i.e. learning on a space of size $\Omega\times S_{n}$, leads to an explosion of the average mean squared error: Learning a function on a fixed number of $k$ objects (say unweighted graphs on $n$ nodes) from $N$ samples leads to a mean squared error of order $\Theta(n!/N)$. In contrast, making use of the inductive bias leads to an average mean squared error of $\Theta(1/N)$. Noting that already on $n=10$ nodes, we have $n!>3\cdot 10^{6}$, we see the clear argument for using the inductive bias.
@@
Figure 2: Struggle of GNN-based and graph-transformer-based models with two WL-equivalent graphs: Graphs A and Graphs B are WL-equivalent, but non-isomorphic. Also physically, they have very different properties, such as different temperature factors (c) and a different cut size. Upon training on a $80\%$ Graph A and $20\%$ Graph B dataset, state-of-the-art GNN-based (EDP-GNN,GDSS) and graph-transformer-based (ConGress) models learn the WL-equivalence class quickly but fail to generate the underlying distribution among the two graphs, with some even predominantly hallucinating WL-equivalent but non-isomorphic graphs (d).
Cost of pushing inductive bias entirely into architecture. One solution would be to impose the inductive bias in the learning architecture. This is what state-of-the-art graph diffusion models do (Niu et al., 2020; Jo et al., 2022; Vignac et al., 2022).
@@
However, using these architectures comes at a cost. As argued in the introduction, since $\mathrm{GI}\in\mathsf{P}$ remains unknown, we expect some limitations. In the case of GNNs this compromise has been precisely characterized: GNNs cannot distinguish between the large families of so-called Weisfeiler-Leman (WL) equivalent graphs (Morris et al., 2019; Xu et al., 2018). For example, all $k$-regular graphs on $n$ vertices for any fixed $k\in\mathbb{N}$ are WL-equivalent (see Lemma 2.1 below and its proof in Appendix L) and therefore indistinguishable for these architectures, which may lead to hallucination.
Lemma 2.1 (WL-equivalence of $k$-regular graphs).
@@
This is a vast class, since e.g. on $n=20$ vertices, there are $510^{\prime}489$ many non-isomorphic, connected 3-regular and thereby WL-equivalent graphs (Meringer, 1999). In particular, we demonstrate on a simple example in Figure 2 that both GNN- and graph-transformer-based methods fail to learn a distribution on two particular WL-equivalent graphs: during sampling, they either fail to learn the distribution on both WL-equivalent graphs or hallucinate other, WL-equivalent but non-isomorphic, graphs (Fig. 2d). This can also be seen during learning: The EDP-GNN model learns the WL-equivalence class quickly (after 500 epochs) but then keeps hallucinating non-isomorphic but WL-equivalent graphs, and in particular struggles to learn the right distribution between graphs $A$ and $B$ for the remaining 4500 epochs (see Appendix O for details). Importantly, those graphs are very different. For instance, if the graphs represented Gaussian Network Models for macromolecules (Tirion, 1996), physical observables such as the temperature factors in X-ray scattering (Haliloglu et al., 1997) would be clearly distinct, see Fig. 2c. Therefore, a diffusion model based on GNNs will suffer from this expressivity blind spot. More generally, any diffusion model relying on a graph-specific learning algorithm will have limited expressivity.

### v2

@@ Abstract
Diffusion models are central to generative modeling and have been adapted to graphs by diffusing adjacency matrix representations. The challenge of having up to $n!$ such representations for graphs with $n$ nodes is only partially mitigated by using permutation-equivariant learning architectures. Despite their computational efficiency, existing graph diffusion models struggle to distinguish certain graph families and their spectra, unless graph data are augmented with ad hoc features. This shortcoming stems from enforcing the inductive bias within the learning architecture. In this work, we leverage random matrix theory to analytically extract the spectral properties of the diffusion process, allowing us to push most of the inductive bias from the architecture into the dynamics. Building on this, we introduce the Dyson Diffusion Model, which employs Dyson’s Brownian motion to capture the spectral dynamics of an Ornstein-Uhlenbeck process on the adjacency matrix. Furthermore, conditioned on the spectral dynamics, we formulate a Lie group diffusion, appropriately modeling the remaining degrees of freedom. Strikingly, the resulting learning problem becomes permutation invariant at the Lie algebra level. We demonstrate that the Dyson Diffusion Model learns graph spectra accurately and outperforms existing graph diffusion models.
@@ 1 Introduction
Diffusion models are a key class of generative models based on noising data with stochastic differential equations (SDEs) and learning their time reversal [1, 2, 3]. They provide state-of-the-art results in many domains such as audio [4] and vision [5]. Generalizing diffusion models from Euclidean space to graphs offers promising applications in numerous areas, such as biology [6] or combinatorial optimization [7]. However, while diffusing adjacency matrix representations is straightforward and popular [8, 9, 10], this approach faces a major obstacle: Each graph with $n$ vertices has up to $n!$ representations as adjacency matrices. Therefore, if one aims to use a diffusion model on the space of matrices, one must learn $n!$ representations per graph. This is not feasible. Previous works tackled this problem by shifting the inductive bias of permutation invariance to the learning algorithm: If the neural network was permutation equivariant, training on one of the (up to $n!$ many) matrix representations would suffice. For example, Niu et al. [8] and Jo et al. [9] use message-passing graph neural networks (GNNs) while ConGress [10] applied graph transformers. However, these learning architectures have a “blind spot” detailed below.
@@
Theoretical Limitations.  The “blind spot” arises from the limited ability of these models to solve Graph Isomorphism ($\mathrm{GI}$): determining if two graphs are structurally identical regardless of node labeling. While permutation equivariance ensures that the model produces consistent outputs for different representations of the same graph, it does not guarantee that the model can differentiate between two structurally different (non-isomorphic) graphs. The failure is a result of how these architectures aggregate information: permutation-symmetric operations – message passing in GNNs and self-attention in graph transformers – can collapse distinct graphs with similar neighborhoods to the same representation, treating them as identical. More formally, $\mathrm{GI}$ is a challenging problem in algorithm theory, and it remains unknown whether $\mathrm{GI}\in\mathsf{P}$ [11]. Since a polynomial time (learning) algorithm perfectly distinguishing all graphs would prove $\mathrm{GI}\in\mathsf{P}$, contemporary (polynomial-time) graph learning algorithms must compromise on expressivity.
@@
Extracting Permutation-Invariant Information from Graph Diffusion. When diffusing an entire adjacency matrix, state-of-the-art work pushes the entire inductive bias into the learning algorithm [8, 9] with possible data augmentation [12, 10, 13]. However, as we show below, this is neither necessary nor desirable (see the previous discussion on blind spots). Using techniques from random matrix theory, we show that an Ornstein–Uhlenbeck (OU) diffusion on the graph can be dissected into diffusion of the (permutation invariant) spectrum and diffusion of the (permutation-dependent) eigenvectors. Our method therefore allows us to learn the spectrum while preserving all remaining information. Moreover, since the spectrum is inherently permutation-invariant, we can parameterize the score using a much broader range of learning architectures, expanding the scope to architectures able to distinguish between graphs that are equivalent in the Weisfeiler-Leman (WL) sense [14, 15]. [^1] We give a theoretical discussion in form of the WL-equivalence class in Section 2 and demonstrate the challenge of those architectures empirically in Figure 2.
@@
Figure 1: Dyson Diffusion Model and its application to graph spectra: A graph on $n$ vertices (a) has up to $n!$ representations as adjacency matrices (b). For an OU-driven diffusion on any adjacency matrix, the permutation-invariant spectrum (c) evolves according to the same SDE (Dyson-BM). A sample path of the $n$ non-intersecting eigenvalues is shown. The marginals of the invariant density for the $\lambda_{k}$ are depicted on the far right. The DyDM diffusion model learns the score $s(\lambda,t)$ (highlighted in yellow) to generate spectra via the time-reversed SDE (4).
@@
Information in the Spectrum. The spectrum of a graph encodes key structural features including connectivity, expansion, and subgraph patterns. Moreover, non-isomorphic WL-equivalent graphs typically have distinct spectra [16]. An idea, therefore, is to augment the graph data based on spectral information [10, 13]. Our work is based on an entirely different method, exploiting analytical insight from random matrix theory to dissect the spectral from the remaining information, allowing to push the inductive bias from the architecture to the dynamics.
@@
Dyson’s Brownian Motion. For an OU process on the space of symmetric matrices, the eigenvalues follow a well-characterized stochastic differential equation, the so-called Dyson Brownian motion (DBM), which is inherently permutation invariant. Thus – in contrast to [8, 9, 10] – the score of DBM can be parameterized with any (not necessarily permutation invariant) neural network. Moreover, contrary to [17], the remaining information is not lost (see Theorem˜3).
Contributions. Our main contributions are:
@@
  • Based on analytical insight, we demonstrate an inherent struggle of GNN-based graph diffusion models.
  • We introduce the Dyson Diffusion Model (DyDM) in Section˜3, extracting the spectral dynamics from an OU-driven diffusion. DyDM learns graph spectra without GNNs or graph transformers; the accompanying conditional eigenvector SDE provides a principled route to modeling the remaining degrees of freedom, which we validate in controlled experiments.
  • We demonstrate in Section˜4 that building on this SDE, DyDM is more effective than existing GNN-based and graph-transformer-based methods for learning graph spectra.
  • We also derive the eigenvector SDE and its time reversal. Expressing the dynamics on the Lie group O(n)O(n)—and its corresponding tangential Lie algebra—enables tractable analysis and efficient simulation. We prove a symmetry exponentially reducing the learning space in the dimension and translate the equivariant learning problem for the eigenvectors into an invariant learning problem on the Lie algebra, opening the field to new architectures.
@@
Beyond graphs, our framework generally applies to symmetric matrices where spectral information is often key. For instance, in statistics, it can encode the importance of principal components [18], while in dynamical systems, [^2] For instance, the generator of Markov dynamics under detailed balance is self-adjoint, see [19]. it reflects the stability and timescales of linear operators.
@@
Notation. We work on the set of symmetric real matrices $\mathrm{Sym}(\mathbb{R}^{n\times n}):=\{A\in\mathbb{R}^{n\times n}:A^{T}=A\}$. The positive integers up to $n$ are denoted by $n]:=\{1,\ldots n\}$. We consider undirected graphs $G=(V,E,w)$ where $V$ is a finite set with edges $E\subseteq\{S\subseteq V:1\leq|S|\leq 2\}$ allowing for self-loops and weights $w:E\to\mathbb{R}$. The family of such graphs of size $n$ is $\mathcal{G}^{n}:=\{G=(V,E,w):|V|=n\}$. For a graph $G\in\mathcal{G}^{n}$ with $\mathbb{R}^{V}=\{f:V\to\mathbb{R}\}$ being the space of functions from $V$ to $\mathbb{R}$, we let $A\equiv A(G):\mathbb{R}^{V}\to\mathbb{R}^{V}$ be the adjacency operator, defined for $f\in\mathbb{R}^{V}$ and $v\in V$ as the weighted sum of $f$ applied to the neighbors of $v$ as $Af(v)\coloneqq\sum_{\{u,v\}\in E}w\left(\{u,v\}\right)f(u).$ As the graphs are undirected, the adjacency operator is self-adjoint with respect to the standard inner product on $\mathbb{R}^{V}$. The operator $A$ therefore has $n$ real eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{n}$. We denote the ordered spectrum of the adjacency operator by $\lambda(G)\coloneqq\lambda(A(G))\coloneqq\{(\lambda_{1},\lambda_{2},\ldots,\lambda_{n})\,:\,\lambda_{1}\geq\ldots\geq\lambda_{n}\}$. Unless defined explicitly otherwise, we use $n\in\mathbb{N}$ for the (vertex) size of the graph, and $N\in\mathbb{N}$ for the number of samples. We denote by $s\sim\mathcal{N}(0,I_{d})$ that the $d$-dimensional random vector $s$ has multivariate normal distribution with 0 mean and unit covariance $I_{d}$.
@@ 2 Limitations of Existing Diffusion Models
We consider the matrix-valued OU-SDE starting from some data $M\mkern-1.0mu(\mkern-1.0mu0\mkern-1.0mu)\!\in\!\mathrm{Sym}\mkern-1.0mu(\mkern-1.0mu\mathbb{R}^{n\times n}\mkern-1.0mu)$ given by
$\mathrm{d}M_{ji}(t)=\mathrm{d}M_{ij}(t)=-\beta M_{ij}(t)\mathrm{d}t+D_{ij}\mathrm{d}B_{ij}(t),\quad 1\!\leq\!i\!\leq\!j\!\leq\!n$ (1)
@@
with $D_{ij}\coloneqq\sqrt{(1+\delta_{ij})\alpha}$ for any constants $\alpha,\beta\in\mathbb{R}^{+}$, where $B_{ij}(t)=B_{ji}(t)$ are independent Brownian motions and $\delta_{ij}=1$ if and only if $i=j$, and 0 otherwise. We consider (1) on the space of symmetric matrices to represent undirected graphs $G\in\mathcal{G}^{n}$. Equation˜1 is an entry-wise OU process that preserves the symmetry of the matrix. Standard diffusion models run (1) until $T>0$ from data samples. To fix notation, denote by $p_{t}$ the distribution of $M(t)$ induced by (1) for $t\geq 0$.
The time-reversal of (1) on time interval $[0,T]$ is a diffusion initialized from $M(T)\!\sim\!p_{T}$} satisfying
$\mathrm{d}M_{ij}(t)=$ $-\left\{\beta M_{ij}(t)+D_{ij}^{2}\left[s(M(t),t)\right]_{ij}\right\}dt+D_{ij}\mathrm{d}\bar{B}_{ij}(t),\quad 1\leq i\leq j\leq n$ (2)
@@
for independent Brownian motions $\bar{B}_{ij}(t)$ [20, 2]. In (2), $s(M,t)$ represents the score matrix at time $t\in(0,T]$, i.e., $s(M,t)\coloneqq\gradient_{M}\ln p_{t}(M)$. This score is intractable but, as the OU process has tractable Gaussian transition densities, we can obtain an estimate $s_{\theta}(M,t)$ of it by minimizing the denoising score matching loss [2]
$L{\mkern-1.0mu}(\theta)\!=\!\mathbb{E}\!\!\left[\!\norm{s_{\theta}(M(t),t)\!\!-\!\!\gradient_{\!M(t)}\!\ln p_{t|0}(M\!(t)|M\!(0))}^{2}_{2}\!\right]\!\!,$ (3)
@@
where the expectation is over $t\sim\mathcal{U}[0,T],M(0)\sim p_{0},M(t)\sim p_{t\mid 0}(\cdot|M(0))$. Samples of $p_{0}$ can then be obtained by simulating an approximation of (2) obtained by sampling $M(T)$ from the invariant distribution of (1), $M_{ij}^{\rm inv}\sim\mathcal{N}(0,\alpha(1+\delta_{ij})/2\beta)$}, and using $s_{\theta}(M,t)$ in place of $s(M,t)$.
@@
Challenges posed by graphs. When working with graphs, to obtain an adjacency matrix for a given weighted graph $G=(V,E,w)\in\mathcal{G}^{n}$, we need to fix an ordering $(v_{1},\ldots,v_{n})$ of all vertices. In fact, given the ordering $(v_{1},\ldots,v_{n})$, the associated matrix $M=(m_{ij})_{1\leq i,j\leq n}\in\mathrm{Sym}(\mathbb{R}^{n\times n})$ has entries $m_{ij}=w(\{v_{i},v_{j}\})$ if $\{v_{i},v_{j}\}\in E$ and 0 otherwise. The challenge is that the adjacency operator $A(G)$ for a given graph $G\in\mathcal{G}^{n}$ admits up to $n!$ distinct adjacency matrices. For instance, in Figure˜1 we see four different matrix representations of the same graph.
@@
Why should we enforce this inductive bias? One could think that the reason for using the inductive bias stems from wanting that graph generative models assign uniform probability to each of the (up to) $n!$ many representations. However, one may easily apply a random independent permutation to the output of the generative model. Instead, the challenge stems from the following problem: In diffusion models, we learn a function (the score) on a set of graphs, say [^3] Importantly, $\Omega$ refers to the set of graphs and not the set of (permutation-sensitive) adjacency matrices. $\Omega$, rather than the distribution directly. Learning on the adjacency representations would correspond to learning on $\Omega\!\times\!S_{n}$. We demonstrate on a toy example in Theorem˜6 that not leveraging the inductive bias, i.e. learning on a space of size $\Omega\!\times\!S_{n}$, leads to an explosion of the average mean squared error; i.e., learning a function on a fixed number of $k$ objects (say unweighted graphs on $n$ nodes) from $N$ samples yields a mean squared error of order $\Theta(n!/N)$. In contrast, using the inductive bias provides an average mean squared error of $\Theta(1/N)$. Given that $n!>3\cdot 10^{6}$ even for n=10n\!=\!10 nodes, utilizing the inductive bias is essential.
@@
Figure 2: Struggle of GNN-based and graph-transformer-based models with two WL-equivalent graphs: Graph A and Graph B are WL-equivalent, but non-isomorphic. Also physically, they have very different properties, such as different temperature factors (c) and a different cut size. Upon training on an $80\%$ Graph A and $20\%$ Graph B dataset, state-of-the-art GNN-based (EDP-GNN, GDSS) and graph-transformer-based (ConGress) models learn the WL-equivalence class quickly but fail to generate the underlying distribution among the two graphs, with some even predominantly hallucinating WL-equivalent but non-isomorphic graphs (d).
@@
Cost of pushing inductive bias entirely into architecture. One solution would be to impose the inductive bias in the learning architecture. This is what state-of-the-art graph diffusion models do [8, 9, 10]. However, using these architectures comes at a cost. As argued in the introduction, since $\mathrm{GI}\in\mathsf{P}$ remains unknown, we expect some limitations. In the case of GNNs this compromise has been precisely characterized: GNNs cannot distinguish between the large families of so-called Weisfeiler-Leman (WL) equivalent graphs [14, 15]. For example, all $k$-regular graphs on $n$ vertices for any fixed $k\in\mathbb{N}$ are WL-equivalent (see Lemma˜1 below and its proof in App. N) and therefore indistinguishable by these architectures. This limitation may lead to hallucination and is also evident in the learned spectra.
Lemma 1 (WL-equivalence of $k$-regular graphs).
@@
This is a vast class, since e.g. on $n=20$ vertices, there are 510,489510{,}489 such non-isomorphic, connected 3-regular and thereby WL-equivalent graphs [21]. In particular, we demonstrate on a simple example in Figure˜2 that both GNN- and graph-transformer-based methods fail to learn a distribution on two particular WL-equivalent graphs: during sampling, they either fail to learn the distribution on both WL-equivalent graphs or hallucinate other, WL-equivalent but non-isomorphic, graphs (Fig. 2d). This can also be observed during learning: The EDP-GNN model learns the WL-equivalence class quickly (after 500 epochs) but then keeps hallucinating non-isomorphic but WL-equivalent graphs, and in particular struggles to learn the right distribution between graphs $A$ and $B$ for the remaining 4500 epochs (see App. R for details). Importantly, those graphs are very different. For instance, if the graphs represented Gaussian Network Models for macromolecules [22], physical observables such as the temperature factors in X-ray scattering [23] would be clearly distinct, see Fig. 2c. Therefore, a diffusion model based on GNNs will suffer from this expressivity blind spot. More generally, any diffusion model relying on a graph-specific learning algorithm will have limited expressivity.